OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

begin by finding the slope of y=4x+7

OpenStudy (anonymous):

a clue please

OpenStudy (anonymous):

look up the slope intercept equation of the line and the point-slope equation of the line.

OpenStudy (anonymous):

slope intercept is y=mx+b

OpenStudy (anonymous):

and point slope equation is y - y1 = m(x - x1)

OpenStudy (anonymous):

So to find the equation of the new line we need a point (which we have) and a slope

OpenStudy (anonymous):

To find the slope of the new line we have to find the slope of the line given in the problem statement. This is because we know that the new line is perpendicular to the old line.

OpenStudy (anonymous):

Therefore, we need to find the slope of the old line from the equation y=4x+7 (which happens to be in slope-intercept form). What is the slope of the line given by y=4x+7?

OpenStudy (anonymous):

y=4x+7 = y-7=4x ?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

for this step we just have to recognize that y=4x+7 is in the same form as y=mx+b

OpenStudy (anonymous):

it is as y=mx+b

OpenStudy (anonymous):

To identify the slope you need to match up what number is in the space where the 'm' is. m in these equations stands for slope.

OpenStudy (anonymous):

oh yea!

OpenStudy (anonymous):

y = 4(2)+7

OpenStudy (anonymous):

y=8+7

OpenStudy (anonymous):

y=15

OpenStudy (anonymous):

no, we don't need to use the point yet.

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

all we are doing first is identifying the slope of the line.

OpenStudy (anonymous):

y=mx+b has a slope of m so y=4x+7 has a slope of ???

OpenStudy (anonymous):

y !

OpenStudy (anonymous):

nope, the slope would be 4 since, four is in the corresponding space as the m in the first equation

OpenStudy (anonymous):

omg :( you are right

OpenStudy (anonymous):

y=mx+b has a slope of m y=4x+b has a slope of 4 Since the four and the m are in the same spots!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

now, do you know the equation to find a perpendicular slope?

OpenStudy (anonymous):

mmm im not sure is it y2-y1/x2-x1 ?

OpenStudy (anonymous):

that is the equation of the slope, given two points.

OpenStudy (anonymous):

right now we are going to calculate the perpendicular slope given any slope.

OpenStudy (anonymous):

Right now we have figured out the slope of the given line, which was 4. The perpendicular slope to 4 is the negative reciprocal of 4.

OpenStudy (anonymous):

\[perpendicular slope=\frac{-1}{m}=\frac{-1}{4}\]

OpenStudy (anonymous):

-1/4=-1/4

OpenStudy (anonymous):

so what do we do now ?

OpenStudy (anonymous):

we now have the slope of the line that we are interested in!

OpenStudy (anonymous):

this means that we have the point and the slope of the new line that we are interested in. This means we can use the point-slope equation of the line.

OpenStudy (anonymous):

yay ! okay so y - y1 = m(x - x1)?

OpenStudy (anonymous):

yes, the m is the slope, y1 is the y point and x1 is the x point from (2, -3)

OpenStudy (anonymous):

2+3 = 4( :s ) ? :0

OpenStudy (anonymous):

what you do for this equation is to plug in m=-1/4; y1=-3; x1=2 leave the rest of the equation the same.

OpenStudy (anonymous):

1+3=-1/4(1-2)

OpenStudy (anonymous):

y + 3 = (-1/4)*(x - 2) do you see how i kept everything the same except for replacing y1, x1, and m?

OpenStudy (anonymous):

yea :0 i though i had to replace it all

OpenStudy (anonymous):

no, when we are looking for the equation of the line, we have to keep x and y.

OpenStudy (anonymous):

Now you need to look up the Standard Form

OpenStudy (anonymous):

:s Ax + By = C ?

OpenStudy (anonymous):

yup, in the standard form A, B, and C have to be integers. Soooooo, the last step that we have to do is change y + 3 = (-1/4)*(x-2) into the form of Ax+By=C

OpenStudy (anonymous):

This is just algebra, begin by multiplying y+3 = (-1/4)*(x-2) by four on both sides.

OpenStudy (anonymous):

y+3= -1/4x - 2.25 ?

OpenStudy (anonymous):

@juantweaver ?